### v2

# An Online Learning Approach to Prompt-based Selection of Generative Models

Abstract
Selecting a sample generation scheme from multiple text-based generative models is typically addressed by choosing the model that maximizes an averaged evaluation score. However, this score-based selection overlooks the possibility that different models achieve the best generation performance for different types of text prompts. An online identification of the best generation model for various input prompts can reduce the costs associated with querying sub-optimal models. In this work, we explore the possibility of varying rankings of text-based generative models for different text prompts and propose an online learning framework to predict the best data generation model for a given input prompt. The proposed PAK-UCB algorithm addresses a contextual bandit (CB) setting with shared context variables across the arms, utilizing the generated data to update kernel-based functions that predict the score of each model available for unseen text prompts. Additionally, we leverage random Fourier features (RFF) to accelerate the online learning process of PAK-UCB and establish a $\widetilde{\mathcal{O}}(\sqrt{T})$ regret bound for the proposed RFF-based CB algorithm over $T$ iterations. Our numerical experiments on real and simulated text-to-image and image-to-text generative models show that RFF-UCB performs successfully in identifying the best generation model across different sample types.
1 Introduction
Text-based generative artificial intelligence (AI) has found numerous applications in various engineering tasks. A prompt-based generative AI represents a conditional generative model that produces samples given an input text prompt. Over the past few years, several frameworks using diffusion models and generative adversarial networks have been proposed to perform text-guided sample generation tasks for various data domains including image, audio, and video (Reed et al.,, 2016; Pan et al.,, 2018; Xu et al.,, 2018; Ding et al.,, 2021; Singer et al.,, 2022; Huang et al.,, 2023; Podell et al.,, 2024). The multiplicity of developed prompt-based models has led to significant interest in developing evaluation mechanisms to rank the existing models and find the best generation scheme. To address this task, several evaluation metrics have been proposed to quantify the fidelity and relevance of samples created by prompt-based generative models, such as CLIPScore (Hessel et al.,, 2021) and PickScore (Kirstain et al.,, 2023).
Example 1 Stable Diffusion v1.5 PixArt-$\bm{\alpha}$
Prompts of Type “dog”
Refer to caption Refer to caption
avg. CLIPScore $36.37\ (\pm 0.13)$ $\bm{37.24}\ (\pm 0.09)$
Prompts of Type “car”
Refer to caption Refer to caption
avg. CLIPScore $\bm{36.10}\ (\pm 0.06)$ $35.68\ (\pm 0.15)$
subfigureExample 1: Stable Diffusion v1.5 attains a higher CLIPScore in generating prompts with term “car” (36.10 versus 35.68) while underperforms for prompts with term “dog” (36.37 versus 37.24).
Example 2 UniDiffuser Stable Diffusion v1.5
Prompts of Type “elephant”
Refer to caption Refer to caption
avg. CLIPScore $\bm{36.67}\ (\pm 0.05)$ $35.08\ (\pm 0.06)$
Prompts of Type “fire hydrant”
Refer to caption Refer to caption
avg. CLIPScore $35.11\ (\pm 0.14)$ $\bm{37.23}\ (\pm 0.05)$
subfigureExample 2: UniDiffuser attains a higher CLIPScore in generating prompts with the term “elephant” (36.67 versus 35.08) while underperforms for prompts with the term “fire hydrant” (35.11 versus 37.23).
Figure 1: Standard text-to-image (T2I) models exhibit different rankings based on CLIPScore (Hessel et al.,, 2021) across the text prompts.
The existing model selection methodologies commonly aim to identify the generative model with the highest relevance score, producing samples that correlate the most with input text prompts. A well-known example is the CLIPScore for image generation models, measuring the expected alignment between the input text and output image of the model using the CLIP embedding (Radford et al., 2021a, ). While the best-model identification strategy has been frequently utilized in generative AI applications, this approach does not consider the possibility that the involved models can perform differently across text prompts. However, it is possible that one model outperforms another model in responding to text prompts from certain categories, while that model performs worse in generating samples for other text categories. Figure 1 shows such examples where two standard text-to-image models exhibit different rankings across the text prompts with terms “dog”/“car” and “elephant”/“fire hydrant”. In general, the different training sets and model architectures utilized to train text-based models can result in the models’ varying performance in response to different text prompts, which is an important consideration in the selection of text-based generative models for various text prompts.
In this work, we aim to develop a learning algorithm to identify the best generative model for a given input prompt, using observed prompt/generated samples collected from the models in the previous sample generation iterations. Since the goal of such text-based model selection is to minimize the data queries from suboptimal generative models for an input text prompt, we view the model selection task as an online learning problem, where after each data generation the learner updates the prediction on which generative model performs the best in response to different text prompts. Here, the goal of the online learner is to utilize the previously generated samples to accurately guess the generation model with the best performance for the incoming text prompt. An optimal online model selection method will result in a bounded regret value, measured in comparison to the sample generation from the groundtruth-best model for the text prompts.
The described online learning task can be viewed as a contextual bandit (CB) problem studied in the bandit literature (Langford and Zhang,, 2007; Li et al.,, 2010). In the CB task, the online learner observes a single context variable (the text prompt in our setting) and predicts the best arm for the current input context. Specifically, we focus on the kernel-based method to predict the score of each available model for an incoming prompt. As the text prompt is shared across the models, the CB task may be suboptimally addressed by the standard linear CB (Chu et al.,, 2011) and kernelized CB (Valko et al.,, 2013) algorithms in the literature, as these algorithms train a single model to predict the reward functions for all the arms. To relax this constraint, we propose Per-Arm Kernelized UCB (PAK-UCB) to address the online prompt-based selection of generative models. According to the PAK-UCB approach, the learner utilizes UCB-scores from arm-specific kernel-based prediction functions to choose the generative model for the incoming text prompt and subsequently update the kernel-based prediction rule based on the generated data for the upcoming iterations. We prove that the proposed PAK-UCB achieves an $\widetilde{O}(\sqrt{T})$ regret bound over a horizon of $T$ iterations.
Since the user applying the CB-based model selection approach may have limited compute power and not be able to afford growing computational costs in the online learning process, we propose to utilize the random Fourier features (RFF) framework (Rahimi and Recht, 2007a, ) to balance the computational load between the iterations of PAK-UCB. We discuss that in the proposed PAK-UCB, the computational cost per iteration will grow cubically as $O(t^{3})$ with iteration $t$. To address the growing cost per iteration of kernel-UCB, we leverage the RFF approach and develop the proxy RFF-UCB algorithm which approximates the solution to PAK-UCB, while the computational costs grow only linearly $O(t)$. We show that the regret bound for PAK-UCB will approximately hold for RFF-UCB, and therefore, RFF-UCB provides an efficient proxy to the PAK-UCB algorithm, which can be run in devices with lower computation budget.
Finally, we present the results of several numerical experiments to show the efficacy of our proposed PAK-UCB and RFF-UCB in the online selection of conditional generative models. In our experiments, we test several pre-trained and simulated text-to-image and image-captioning (image-to-text) models, where different models lead to different rankings of CLIPScore values across sample types. Our numerical results suggest a fast convergence of the proposed online learning algorithms to the best model available for different prompt types. Moreover, we apply the RFF-UCB method and show how the algorithm can infer the model with higher CLIPScore with a growing accuracy as the iteration grows. In our experiments, the proposed PAK-UCB and RFF-UCB outperform various baseline methods including the standard LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) methods. The main contributions of our work can be summarized as:
  • Studying the prompt-based selection of conditional generative models to improve the performance scores over every individual model
  • Developing the contextual bandit-based PAK-UCB and RFF-UCB algorithms for the online selection of prompt-based generative models
  • Providing the theoretical analysis of the regret and computational costs of PAK-UCB and RFF-UCB online learning methods
  • Presenting numerical results on the online selection of generative models based on the incoming prompt using PAK-UCB and RFF-UCB
2 Related Work
(Automatic) Evaluation of conditional generative models. Evaluating the conditional generative models has been studied extensively in the literature. For text-to-image (T2I) generation, earlier methods primarily rely on the Inception score (Salimans et al.,, 2016) and Fréchet inception distance (Heusel et al.,, 2017). More recent works propose reference-free metrics for robust automatic evaluation of T2I and image captioning, with notable examples being CLIPScore (Hessel et al.,, 2021) and PickScore (Kirstain et al.,, 2023). Kim et al., (2022) propose a mutual-information-based metric, which attains consistency across benchmarks, sample parsimony, and robustness. To provide a holistic evaluation of T2I models, several works focus on multi-objective evaluation. Astolfi et al., (2024) propose to evaluate conditional image generation in terms of prompt-sample consistency, sample diversity, and fidelity. Kannen et al., (2024) introduce a framework to evaluate T2I models regarding cultural awareness and cultural diversity. Masrourisaadat et al., (2024) examine the performance of several T2I models in generating images such as human faces and groups and present a social bias analysis. Another line of study explores evaluation approaches using large language models (LLMs). Tan et al., (2024) develop LLM-based evaluation protocols that focus on the faithfulness and text-image alignment. Peng et al., (2024) introduce a GPT-based benchmark for evaluating personalized image generation. For evaluation of text-to-video (T2V) generation, Huang et al., (2024) introduce VBench as a comprehensive evaluation of T2V models in terms of quality and consistencey.
(Kernelized) Contextual bandits.  The contextual bandits (CB) is an efficient framework for online decision-making with side information (Langford and Zhang,, 2007; Foster et al.,, 2018), which is widely adopted in domains such as recommendation system and online advertisement (Li et al.,, 2010). A key to its formulation is the relationship between the context (vector) and the expected reward. In linear CB, the reward is assumed to be linear to the context vector (Li et al.,, 2010; Chu et al.,, 2011). To incorporate non-linearity, Valko et al., (2013) propose kernelized CB, which assumes the rewards are linear-realizable in a reproducing kernel Hilbert space (RKHS). To address the growing computation, recent work leverages the assumption that the kernel matrix is often approximately low-rank and uses Nyström approximations (Calandriello et al.,, 2019, 2020; Zenati et al.,, 2022).
3 Preliminaries
3.1 CLIPScore
CLIPScore (Hessel et al.,, 2021) is a widely used automatic metric to evaluate the alignment of text-to-image/video (T2I/V) and image captioning models. Let $(y,x)\in{\mathcal{Y}}\times{\mathcal{X}}$ be any text-image pair. We denote by $\bm{c}_{y}\in{\mathbb{S}}^{d-1}:=\{z\in{\mathbb{R}}^{d}:\|z\|_{2}=1\}$ and $\bm{v}_{x}\in{\mathbb{S}}^{d-1}$ the (normalized) embeddings of text $y\in{\mathcal{Y}}$ and image $x\in{\mathcal{X}}$, respectively, both extracted by CLIP (Radford et al., 2021b, ). The CLIPScore (Hessel et al.,, 2021) is given by
$\textup{CLIPScore}^{\textup{T2I}}(y,x):=\max\{0,100\cdot\cos(\bm{v}_{x},\bm{c}_{y})\},$ (1)
and note that $\cos(\bm{v}_{x},\bm{c}_{y})=\langle\bm{v}_{x},\bm{c}_{y}\rangle$ as we operate under the normalized embeddings. Further, for a video $X:=\{x^{(l)}\}_{l=1}^{L}$ consisting of $L$ frames, where $x^{(l)}$ is the $l$-th frame, the score is the averaged frame CLIPScore, that is,
$\textup{CLIPScore}^{\textup{T2V}}(y,X):=\frac{1}{L}\sum_{l=1}^{L}\textup{CLIPScore}^{\textup{T2I}}(y,x^{(l)}).$ (2)
3.2 Kernel Methods and Random Fourier Features
Let $\phi:{\mathbb{R}}^{d}\to{\mathcal{H}}$ denote a mapping from the primal space ${\mathbb{R}}^{d}$ to the (possibly infinite-dimensional) associated reproducing kernel Hilbert space (RKHS) ${\mathcal{H}}$. The corresponding kernel function is defined by $k(y,y^{\prime}):=(\phi(y))^{\top}\phi(y^{\prime})$ for any $y,y^{\prime}\in{\mathbb{R}}^{d}$. The kernel function $k$ is positive definite if $\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}k(y_{i},y_{j})\geq 0$ for any $n\in{\mathbb{N}}_{+}$, $y_{1},\cdots,y_{n}\in{\mathbb{R}}^{d}$, and $c_{1},\cdots,c_{n}\in{\mathbb{R}}$. Further, a positive definite kernel function is shift invariant if $k(y,y^{\prime}):=k(y-y^{\prime})$ for any $y,y^{\prime}\in{\mathbb{R}}^{d}$. An example is the radial basis function (RBF) kernel:
$k_{\textup{RBF}}(y,y^{\prime})=\exp(-\|y-y^{\prime}\|_{2}^{2}/(2\sigma^{2})),$
where $\sigma>0$.
Kernel ridge regression (KRR). Given empirical data $(y_{1},s_{1}),\cdots,(y_{n},s_{n})$, where $\{y_{i}\in{\mathbb{R}}^{d}\}_{i=1}^{n}$ are dependent variables and $\{s_{i}\in{\mathbb{R}}\}_{i=1}^{n}$ are target variables, respectively, the kernel method assumes the existence of $w^{\star}\in{\mathcal{H}}$ such that $\mathbb{E}[s_{i}|y_{i}]=(\phi(y))^{\top}w^{\star}$ for any $i=1,\cdots,n$. Let $\alpha\geq 0$ denote a regularization parameter. KRR constructs the estimator
$\widehat{s}_{\textup{KRR}}(y):=k_{y}^{\top}(K+\alpha I_{n})^{-1}v$
for any $y\in{\mathbb{R}}^{d}$, where $K=[k(y_{i},y_{j})]_{i,j=1}^{n}\in{\mathbb{R}}^{n\times n}$ is the kernel matrix, $v:=[s_{1},\cdots,s_{n}]^{\top}\in{\mathbb{R}}^{n}$, and $k_{y}=[k(y_{1},y),\cdots,k(y_{n},y)]^{\top}\in{\mathbb{R}}^{n}$. The estimator can be interpreted as first estimating $w^{\star}$ by ridge regression
$\widehat{w}:=\operatorname*{arg\,min}_{w\in{\mathcal{H}}}\sum_{i=1}^{n}((\phi(y_{i}))^{\top}w-s_{i})^{2}+\alpha\|w\|,$
where $\|w\|:=\sqrt{w^{\top}w}$ for any $w\in{\mathcal{H}}$, and then making the prediction $\widehat{s}_{\textup{KRR}}(y)=(\phi(y))^{\top}\widehat{w}$.
Random Fourier features (RFF). One problem of KRR is that it scales poorly with the size $n$ of the empirical data, i.e., computing the KRR estimator generally requires $\Omega(n^{3})$ time and $\Omega(n^{2})$ memory. To address this problem, Rahimi and Recht, 2007b propose to scale up kernel methods by RFF sampling. Specifically, the Bochner’s Theorem (Rudin,, 2017) implies that for any (properly scaled) shift-invariant kernel $k(y,y^{\prime})=k(y-y^{\prime})$, there exists a distribution $p\in\Delta({\mathbb{R}}^{d})$ such that $k(y,y^{\prime})=\mathbb{E}_{w\sim p}[e^{iw^{\top}(y-y^{\prime})}],$ where $e^{i\theta}:=\cos\theta+i\cdot\sin\theta$ for any $\theta\in{\mathbb{R}}$ and $i$ is the imaginary unit. Therefore, the idea of RFF is to sample $w_{1},\cdots,w_{s}\sim p$ and approximate $k(y,y^{\prime})$ by the empirical mean $s^{-1}\sum_{j=1}^{s}e^{iw_{j}^{\top}(y-y^{\prime})}$ to within $\epsilon$ with only $s=O(d\epsilon^{-2}\log(1/\epsilon^{2}))$. Since the kernel $k(\cdot)$ is real, we can replace the complex exponentials with cosines and define
$\varphi(y):=\sqrt{\frac{2}{s}}\cdot[\cos(w_{1}^{\top}y+b_{1}),\cdots,\cos(w_{s}^{\top}y+b_{s})]^{\top},$ (3)
[…]
$\widetilde{s}_{\textup{KRR}}(y):=(\widetilde{\Phi}^{*}\widetilde{\Phi}+\alpha I_{s})^{-1}\widetilde{\Phi}^{*}v,$
where $\Phi:=[\varphi(y_{i})^{\top}]_{i=1}^{n}\in{\mathbb{C}}^{n\times s}$, can be computed in $O(ns^{2})$ time and $O(ns)$ memory, giving substantial computational savings if $s\ll n$ (Avron et al.,, 2017). For the RBF kernel, the distribution $p_{\textup{RBF}}$ is the multivariate Gaussian ${\mathcal{N}}(0,\sigma^{-2}\cdot I_{d})$.
4 Prompt-Based Selection as Contextual Bandits
In this section, we introduce the framework of online prompt-based selection of generative models, which is given in Protocol 1. Let $[N]:=\{1,\cdots,N\}$ for any positive integer $N\in{\mathbb{N}}_{+}$. We denote by ${\mathcal{G}}:=[G]$ the set of (prompt-based) generative models. The evaluation proceeds in $T\in{\mathbb{N}}_{+}$ iterations.
[…]
$\textup{Regret}(T):=\sum_{t=1}^{T}\left(s_{\star}(y_{t})-s_{g_{t}}(y_{t})\right),$ (4)
where we denote by $s_{g}(y):=\mathbb{E}_{x_{g}\sim P_{g}(\cdot|y)}[s(y,x_{g})]$ the expected score of any model $g\in{\mathcal{G}}$ and $s_{\star}(y):=\max_{g\in{\mathcal{G}}}s_{g}(y)$ the optimal expected score, both conditioned on the prompt $y$.
Protocol 1 Online Prompt-based Selection of Generative Models
[…]
2:observation sequence ${\mathcal{D}}\leftarrow\varnothing$
3:for iteration $t=1,2,\cdots,T$ do
4:     Prompt $y_{t}\sim\rho$ is revealed.
5:     Algorithm ${\mathcal{A}}$ picks model $g_{t}\sim{\mathcal{A}}(\cdot|{\mathcal{D}},y_{t})$ and samples an answer $x_{t}\sim P_{g_{t}}(\cdot|y_{t})$.
6:     Score $s_{t}\leftarrow s(y_{t},x_{t})$ is assigned.
7:     Update observation ${\mathcal{D}}\leftarrow{\mathcal{D}}\cup\{(y_{t},g_{t},s_{t})\}$.
8:end for
5 An Optimism-based Approach for Prompt-based Selection
Under the online prompt-based selection setting, a key challenge is to learn the relationship between the prompt and the score of each model. In this paper, we consider kernel methods for score prediction. Specifically, for each model $g\in{\mathcal{G}}$, we assume the existence of a (possibly non-linear and infinite-dimensional) feature map $\phi$ and a weight $w_{g}^{\star}$ in an RKHS, such that the score $s_{g}(y)$ conditioned on the prompt $y$ is given by $(\phi(y))^{\top}w_{g}^{\star}$. Notably, the weight vector $w_{g}^{\star}$ is arm-specific and can vary across the models, which is stated in the following assumption.
Assumption 1 (Realizability).
There exists a mapping $\phi:{\mathbb{R}}^{d}\to{\mathcal{H}}$ and weight $w_{g}^{\star}\in{\mathcal{H}}$ such that score $s_{g}(y)=\langle y,w_{g}^{\star}\rangle_{\mathcal{H}}$ for any prompt vector $y\in{\mathbb{R}}^{d}$ and model $g\in{\mathcal{G}}$. Further, it holds that $\|w_{g}^{\star}\|\leq 1$, and $k(y,y)\leq\kappa^{2}$ and $\|\phi(y)\|\leq 1$ for any $y\in{\mathcal{Y}}$, where $k:{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ is the kernel function of the mapping $\phi$.
Remark 1 (Comparison to linear and kernelized CB).
We note that Assumption 1 differs from the standard settings of linear and kernelized contextual bandit (CB) (Chu et al.,, 2011; Valko et al.,, 2013; Zenati et al.,, 2022). Specifically, in these settings, a different context variable is observed at each arm, and a single weight vector is applied uniformly across all arms. In the prompt-based model selection setting, the text prompt (i.e., the context variable) is shared among arms (i.e., the generative models) and can result in different performance scores across the models, which cannot be captured by the standard kernelized CB approach.
Remark 2 (Kenelized scores).
Note that the relationship between the prompt vector and score can be non-linear and dependent on the choice of the generative model. For instance, the generated images often vary across the prompts and different T2I models, which will affect the resulting CLIPScore (1).
5.1 The PAK-UCB Algorithm
In this section, we present PAK-UCB in Algorithm 2, an online learning approach to prompt-based model selection. For each incoming prompt, PAK-UCB first estimates the performance scores via kernel ridge regression (KRR) and then picks the model with the highest estimated score. Unlike standard LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) algorithms, which train a single reward function shared across all arms, PAK-UCB learns arm-specific functions to predict the score of each model.
[…]
2:observation sequence ${\mathcal{D}}\leftarrow\varnothing$ and index set $\Psi_{g}\leftarrow\varnothing$ for all $g\in{\mathcal{G}}$
3:for iteration $t=1,2,\cdots,T$ do
4:     Prompt $y_{t}\sim\rho$ is revealed.
5:     Compute $(\widehat{\mu}_{g},\widehat{\sigma}_{g})\leftarrow\textsc{Compute\_UCB}({\mathcal{D}},y_{t},\Psi_{g})$ for each $g\in{\mathcal{G}}$.
6:     Pick model $g_{t}\leftarrow\operatorname*{arg\,max}_{g\in{\mathcal{G}}}\{\widehat{s}_{g}\}$, where $\widehat{s}_{g}\leftarrow\widehat{\mu}_{g}+(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{g}$.
7:     Sample an answer $x_{t}\sim P_{g_{t}}(\cdot|y_{t})$ and compute the score $s_{t}\leftarrow s(y_{t},x_{t})$.
8:     Update ${\mathcal{D}}\leftarrow{\mathcal{D}}\cup\{(y_{t},s_{t})\}$ and $\Psi_{g_{t}}\leftarrow\Psi_{g_{t}}\cup\{t\}$.
9:end for
10:
11:function Compute_UCB(${\mathcal{D}},y,\Psi_{g}$)
12:     if $\Psi_{g}$ is empty then
13:         $\widehat{\mu}_{g}\leftarrow+\infty,\widehat{\sigma}_{g}\leftarrow+\infty$.
14:     else
15:         Set $K\leftarrow[k(y_{i},y_{j})]_{i,j\in\Psi_{g}},v\leftarrow[s_{i}]_{i\in\Psi_{g}}^{\top}$, and $k_{y}\leftarrow[k(y,y_{i})]_{i\in\Psi_{g}}^{\top}$.
16:         $\widehat{\mu}_{g}\leftarrow k_{y}^{\top}(K+\alpha I)^{-1}v$.
17:         $\widehat{\sigma}_{g}\leftarrow\alpha^{-\frac{1}{2}}\sqrt{k(y,y)-k_{y}^{\top}(K+\alpha I)^{-1}k_{y}}$.
18:     end if
19:     return $(\widehat{\mu}_{g},\widehat{\sigma}_{g})$.
20:end function
Algorithm 2 Per-Arm Kernelized UCB (PAK-UCB)
The key component in PAK-UCB is the function Compute_UCB (lines 8-17), which outputs both the KRR estimator $\widehat{\mu}_{g}$ and an uncertainty quantifier $\widehat{\sigma}_{g}$. As the weight vector $w_{g}^{\star}$ can vary across the arms, PAK-UCB constructs the KRR dataset using prompt-score pairs from iterations where model $g$ is chosen, with the corresponding indices stored in the set $\Psi_{g}$ (line 6). The estimated score is then computed by $\widehat{s}_{g}=\widehat{\mu}_{g}+(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g}$ (line 4), which is initially set to +$\infty$+ ∞ to ensure each model is picked at least once (lines 9-10). Particularly, the following lemma shows that under some conditions, $\widehat{s}_{g}$ is an optimistic estimation of $s_{g}(y_{t})$ with high probability. The detailed proof can be found in Appendix C.1.
Lemma 1 (Optimism).
Let $\Psi_{g}\subseteq[T]$ be an index set such that the set of scores $\{s_{t}:t\in\Psi_{g}\}$ are independent random variables. Then, under Assumption 1, with probability at least $1-\delta$, the quantity $\widehat{\mu}_{g}$ computed in function Compute_UCB$({\mathcal{D}},y,\Psi_{g})$ satisfies that
$\left|\widehat{\mu}_{g}-s_{g}(y)\right|\leq(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g},$ (5)
where $\eta=\sqrt{2\log(2/\delta)}$. Hence, it holds that $\widehat{s}_{g}=\widehat{\mu}_{g}+(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g}\geq s_{g}(y)$.
We show that a variant of PAK-UCB attains a regret of $\widetilde{O}(\sqrt{GT})$. The formal statement and the proof can be found in Appendix A.
Theorem 1 (Regret, informal).
Under the same conditions in Lemma 1, with probability of at least $1-\delta$, a variant of Algorithm 2 attains a regret of $\widetilde{O}(\sqrt{GT})$.
5.2 PAK-UCB with Random Fourier Features
The PAK-UCB solves a KRR for each model at an iteration to estimate the scores, which can be expensive in both computation and memory for a large number of iterations. To address this problem, we leverage the random Fourier features (RFF) sampling (Rahimi and Recht, 2007b, ) for positive definite shift-invariant kernels. At a high level, RFF maps the input data, e.g., the prompt (vector) in our setting, to a randomized low-dimensional feature space and then applies fast linear methods to solve the regression problem. Particularly, the inner product between these projected randomized features is an unbiased estimation of the kernel value.
We present the RFF-UCB algorithm, which is a variant of PAK-UCB with random features. RFF-UCB leverages an RFF-based approach to compute the mean and uncertainty quantifier in line 3 of Algorithm 2, which we present in Algorithm 3. Upon receiving the regression dataset consisting of prompt-score pairs, Compute_UCB_RFF first projects each $d$-dimensional prompt vector to a randomized $s$-dimensional feature space according to Equation (3) (lines 5-6) and then solves a linear ridge regression to estimate the mean and uncertainty (lines 8-9). To derive statistical guarantees, the number of features varies according to both the input data and error thresholds, which we will specify in Appendix B.1. In practice, we find that a size around 50 can attain satisfactory empirical performance. To see why RFF can reduce the computation, note that the size of the (regularized) Gram matrix $(\widetilde{\Phi}_{g}^{\top}\widetilde{\Phi}_{g}+\alpha I)$ in line 8 is fixed to be $s$ in the whole process, while the size of $(K+\alpha I)$ in line 13 of Algorithm 2 scales with $|\Psi_{g}|$ and can grow linearly over iterations. Particularly, the following lemma shows that Compute_UCB_RFF can reduce the time and space by an order of $O(t^{2})$ and $O(t)$, respectively.
1:the Fourier transform $p$ of a positive definite shift-invariant kernel $k(y,y^{\prime})=k(y-y^{\prime})$, error thresholds $\operatorname{\epsilon_{\textup{RFF}}},\operatorname{\Delta_{\textup{RFF}}}>0$, regularization and exploration parameters $\alpha,\eta\geq 0$
2:number of features $s$, bonus terms ${\mathcal{B}}_{g,1}$ and ${\mathcal{B}}_{g,2}$
3:function Compute_UCB_RFF(${\mathcal{D}},y,\Psi_{g}$)
4:     if $\Psi_{g}$ is empty then
5:         $\widetilde{\mu}_{g}\leftarrow+\infty,\widetilde{\sigma}_{g}\leftarrow+\infty$.
6:     else
[…]
8:         Define mapping $\varphi(y^{\prime})\leftarrow\sqrt{\frac{2}{s}}\cdot[\cos(w_{1}^{\top}y+b_{1}),\cdots,\cos(w_{s}^{\top}y+b_{s})]^{\top}$ for any $y^{\prime}\in{\mathbb{R}}^{d}$.
9:         Set $\widetilde{\Phi}_{g}\leftarrow[\varphi(y_{i})^{\top}]_{i\in\Psi_{g}}$ and $v\leftarrow[s_{i}]_{i\in\Psi_{g}}^{\top}$.
10:         $\widetilde{\mu}_{g}\leftarrow(\varphi(y))^{\top}(\widetilde{\Phi}_{g}^{\top}\widetilde{\Phi}_{g}+\alpha I)^{-1}\widetilde{\Phi}_{g}^{\top}v+{\mathcal{B}}_{g,1}$.
11:         $\widetilde{\sigma}_{g}\leftarrow\alpha^{-\frac{1}{2}}(1-(\varphi(y))^{\top}(\widetilde{\Phi}_{g}^{\top}\widetilde{\Phi}_{g}+\alpha I)^{-1}\widetilde{\Phi}_{g}^{\top}\widetilde{\Phi}_{g}(\varphi(y)))^{\frac{1}{2}}+{\mathcal{B}}_{g,2}$.
12:     end if
13:     return $(\widetilde{\mu}_{g},\widetilde{\sigma}_{g})$.
14:end function
Algorithm 3 Compute UCB with Random Fourier Features
Lemma 2 (Time and space complexity).
At any iteration $t\in[T]$, Compute_UCB (lines 8-17 of Algorithm 2) requires $O(t^{3}/G^{2})$ time and $O(t^{2}/G)$ space, while Compute_UCB_RFF with random features of size $s\in{\mathbb{N}}_{+}$ (Algorithm 3) requires $O(ts^{2})$ time and $O(ts)$ space, where $G$ is the number of generators. See Appendix B.5 for details.
It can be shown that the implementation of PAK-UCB with RFF attains the exact same regret guarantees for adaptively selected feature sizes. The formal statement and the proof can be found in Appendix B.2.
Theorem 2 (Regret when using RFF, informal).
Under the same conditions in Theorem 4, a variant of RFF-UCB attains a regret of $\widetilde{O}(\sqrt{GT})$.
6 Numerical Results
In this section, we present numerical results for the proposed 1) PAK-UCB-poly3 algorithm (PAK-UCB using a polynomial kernel with degree 3, i.e., $k_{\text{poly3}}(x_{1},x_{2})=(1+x_{1}^{\top}x_{2})^{3}$), and 2) the RFF-UCB algorithm with the RBF kernel. Our primary focus is on prompt-based selection among standard text-to-image (T2I) models: Stable Diffusion v1.5111https://huggingface.co/docs/diffusers/en/api/pipelines/stable_diffusion/text2img, PixArt-$\alpha$-XL-2-512x512222https://huggingface.co/PixArt-alpha/PixArt-XL-2-512x512, UniDiffuser [^3] https://github.com/thu-ml/unidiffuser, and DeepFloyd IF-I-M-v1.0.444https://github.com/deep-floyd/IF. Additionally, we conduct experiments on image captioning (image-to-text) task and video data under synthetic setups. Detailed implementation of the algorithms and additional results can be found in Appendix D.
6.1 Experimental Details
Baselines. We compare PAK-UCB-poly3 and RFF-UCB with five baselines, including 1) LinUCB and KernelUCB-poly3: standard LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) using a polynomial kernel with degree 3, where the context variable of each arm is given by the concatenation of the CLIP embedding of the text prompt and its one-hot indicator vector, 2) One-arm Oracle: always picking the model with the maximum averaged CLIPScore, 3) PAK-UCB-lin: PAK-UCB with linear kernel, i.e., $k_{\text{lin}}(x_{1},x_{2})=x_{1}^{\top}x_{2}$, which does not incorporate non-linearity in score estimation, and 4) Naive-KRR: PAK-UCB-poly3 without exploration, which selects the model with the highest estimated mean conditioned to the prompt.
Performance metrics. For each experiment, we report two performance metrics: (i) outscore-the-best (O2B): the difference between the CLIPScore attained by the algorithm and the highest average CLIPScore attained by any single model, and (ii) optimal-pick-ratio (OPR): the overall ratio that the algorithm picks the best generator conditioned to the prompt type.
6.2 Summary
The main finding of our numerical experiments is the improvement of the proposed contextual bandit PAK-UCB algorithm over the one-arm oracle baseline. This result means that the online learning algorithm can outperform a user with side-knowledge of the single best-performing model, which is made possible by a prompt-based selection of the model. This finding supports the application of contextual bandit algorithms in the selection of text-based generative models. Moreover, our numerical results indicate that the proposed PAK-UCB algorithm can significantly outperform the standard LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) methods, showcasing the effectiveness of our framework in the prompt-based model selection setting. Additionally, the proposed PAK-UCB algorithm can perform better with a non-linear kernel function. Finally, in our experiments, the proposed RFF-UCB variant could reduce the computational costs of the general PAK-UCB algorithm.
6.3 Experimental Setups
Setup 1. Prompt-based selection among standard text-to-image models. The first set of experiments are on the setup illustrated in Figure 1(a), where we generate images from Stable Diffusion v1.5 and PixArt-$\alpha$. The results show that PAK-UCB-poly3 outperforms the baseline algorithm and attains a high optimal-pick-ratio, which shows that it can identify the optimal model conditioned to the prompt (see Figure LABEL:fig:t2itab1). Additionally, we provide numerical results on various T2I generative models, including UniDiffuser and DeepFloyd (see Figure LABEL:fig:t2itab2).
Setup 2. Adaptation to new prompt types and models. We consider scenarios where new text-to-image models or prompt types are introduced after the initial deployment. At the beginning of the first experiment, Stable Diffusion v1.5 and PixArt-$\alpha$ are available, and UniDiffuser is introduced after $2500$ iterations. In the second experiment, we generate samples from both PixArt-$\alpha$ and UniDiffuser, and a new prompt type is introduced after each $1000$ iterations. The results show that PAK-UCB-poly3 can well adapt to new prompt types and generators (see Figure LABEL:fig:varying_model).
Setup 3. Synthetic experiments on other conditional generation tasks. We provide numerical results on image-captioning (image-to-text) and text-to-video (T2V) task under synthetic setups. The details can be found in Appendix D.
7 Conclusion
In this work, we investigated prompt-based selection of generative models using a contextual bandit algorithm, which can identify the best available generative model for a given text prompt. We propose two kernel-based algorithms, including PAK-UCB and RFF-UCB, to perform this selection task. Our numerical results on text-to-image, text-to-video, and image-captioning tasks demonstrate the effectiveness of the proposed framework in scenarios where the available generative models have varying performance rankings depending on the type of prompt. An interesting direction for future research is to extend the application of our algorithms to text-to-text language models, where different models may respond better to questions on different topics. Furthermore, considering evaluation criteria beyond relevance, such as diversity and novelty scores, could lead to extensions of our proposed framework.
Appendix A Proof in Section 5.1
The technical challenge in analyzing PAK-UCB is that predictions in later iterations make use of previous outcomes. Hence, the rewards $\{s_{i}\}_{i\in\Psi_{g}}$ are not independent if the index set $\Phi_{g}$ is updated each time when model $g$ is chosen (line 6 of Algorithm 2). To address this problem, we leverage a standard approach used in prior works (Auer,, 2003; Chu et al.,, 2011; Valko et al.,, 2013) and present a variant of PAK-UCB in Algorithm 4, which is called Sup-PAK-UCB. We prove it attains a regret of order $\widetilde{O}(\sqrt{T})$.
A.1 The Sup-PAK-UCB Algorithm
[…]
2:observation sequence ${\mathcal{D}}\leftarrow\varnothing$ and index sets $\{\Psi_{g}^{m}\leftarrow\varnothing\}_{m=1}^{M}$ for all $g\in{\mathcal{G}}$, where $M\leftarrow\log T$
3:for iteration $t=1,2,\cdots,T$ do
4:     Prompt $y_{t}\sim\rho$ is revealed.
5:     Set stage $m\leftarrow 1$ and $\widehat{{\mathcal{G}}}^{1}\leftarrow{\mathcal{G}}$.
6:     repeat
7:         Compute $\{(\widehat{\mu}_{g}^{m},\widehat{\sigma}_{g}^{m})\leftarrow\textsc{Compute\_UCB}({\mathcal{D}},y_{t},\Psi_{g}^{m})\}_{g\in\widehat{{\mathcal{G}}}^{m}}$.
8:         Set upper confidence bound $\widehat{s}_{g}^{m}(y)\leftarrow\widehat{\mu}_{g}^{m}+(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{g}^{m}$ for all $g\in\widehat{{\mathcal{G}}}^{m}$.
9:         if $(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{g}^{m}\leq 1/\sqrt{T}$ for all $g\in\widehat{{\mathcal{G}}}^{m}$ then
10:              Pick model $g_{t}\leftarrow\operatorname*{arg\,max}\widehat{s}_{g}^{m}(y)$.
11:         else if $(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{g}^{m}\leq 2^{-m}$ for all $g\in\widehat{{\mathcal{G}}}^{m}$ then
12:              $\widehat{{\mathcal{G}}}^{m+1}\leftarrow\{g\in\widehat{{\mathcal{G}}}^{m}:\widehat{s}_{g}^{m}(y)\geq\max_{g\in\widehat{{\mathcal{G}}}^{m}}\widehat{s}_{g}^{m}(y)-2^{1-m}\}$.
13:              Set stage $m\leftarrow m+1$.
14:         else
15:              Pick $g_{t}\in\widehat{{\mathcal{G}}}^{m}$ such that $(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{g}^{m}>2^{-m}$.
16:              Update $\Psi_{g_{t}}^{m}\leftarrow\Psi_{g_{t}}^{m}\cup\{t\}$.
17:         end if
18:     until a model $g_{t}$ is selected
19:     Sample an answer $x_{t}\sim P_{g_{t}}(\cdot|y_{t})$ and compute the score $s_{t}\leftarrow s(y_{t},x_{t})$.
20:     Update ${\mathcal{D}}\leftarrow{\mathcal{D}}\cup\{(y_{t},s_{t})\}$.
21:end for
Algorithm 4 Sup-PAK-UCB
A.2 Analysis
In this section, we prove the following regret bound of Sup-PAK-UCB.
Theorem 3 (Regret of Sup-PAK-UCB).
Under Assumption 1, with probability at least $1-\delta$, the regret of Sup-PAK-UCB with $\eta=\sqrt{2\log(2(\log T)GT/\delta)}$ is bounded by
$\textup{Regret}(T)\leq\widetilde{O}\left((1+\sqrt{\alpha})\sqrt{d_{\textup{eff}}\left(10+\frac{15}{\alpha}\right)GT}\right)$ (6)
where $d_{\textup{eff}}$ is a data-dependent quantity defined in Lemma 10 and logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$.
Notations. To facilitate the analysis, we add an subscript $t$ to all the notations in Algorithm 4 to indicate they are quantities computed at the $t$-th iteration, i.e., $\widehat{\mu}_{g,t}^{m},\widehat{\sigma}_{g,t}^{m},\widehat{s}_{g,t}^{m},\widehat{{\mathcal{G}}}_{t}^{m}$, and $\Psi_{g,t}^{m}$. [^5] Note that line 14 in the algorithm is rewritten as “$\Psi_{g_{t},t+1}^{m}\leftarrow\Psi_{g_{t},t+1}^{m}\cup\{t\},\Psi_{g_{t},t+1}^{m^{\prime}}\leftarrow\Psi_{g_{t},t}^{m^{\prime}}$ for any $m^{\prime}\neq m$, and $\Psi_{g,t+1}^{m^{\prime}}\leftarrow\Psi_{g,t}^{m^{\prime}}$ for any $g\neq g_{t}$ and $m\in[M]$”. In addition, we set $\Psi_{g,t+1}^{m}\leftarrow\Psi_{g,t+1}^{m}$ for all $g\in{\mathcal{G}}$ and $m\in[M]$ in line 8.
We leverage the following two lemmas to prove Theorem 3. The first lemma shows that the construction of index sets $\{\Psi_{g,t}^{m}\}_{m=1}^{M}$ ensures the independence among the rewards $\{s_{i}\}_{i\in\Psi_{g,t}^{m}}$, which allows us to utilize Lemma 1 to bound the estimation error.
Lemma 3 (Auer, (2003), Lemma 14).
For any iteration $t\in[T]$, model $g\in{\mathcal{G}}$, and stage $m\in[M]$, the set of rewards $\{s_{i}\}_{i\in\Psi_{g,t}^{m}}$ are independent random variables such that $\mathbb{E}[s_{i}]=s_{g}(y_{i})$.
The second lemma shows several properties of the estimated score $\widehat{s}_{g,t}^{m}$ and the set $\widehat{{\mathcal{G}}}^{m}_{t}$. The detailed proof can be found in Appendix C.2.
Lemma 4 (Valko et al., (2013), Lemma 7).
With probability at least $1-(MGT)\delta$, for any iteration $t\in[T]$ and stage $m\in[M]$, the following hold:
  • $|\widehat{\mu}_{g,t}^{m}-s_{g}(y_{t})|\leq(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g,t}^{m}$ for any $g\in\widehat{{\mathcal{G}}}^{m}_{t}$,
  • $\operatorname*{arg\,max}_{g\in{\mathcal{G}}}s_{g}(y_{t})\in\widehat{{\mathcal{G}}}^{m}_{t}$, and
  • $s_{\star}(y_{t})-s_{g}(y_{t})\leq 2^{3-m}$ for any $g\in\widehat{{\mathcal{G}}}^{m}_{t}$.
Now, we are ready to finish the proof of Theorem 3.
Proof of Theorem 3.
Let ${\mathcal{T}}_{1}:=\cup_{m\in[M],g\in{\mathcal{G}}}\Psi_{g,T+1}^{m}$ and ${\mathcal{T}}_{0}:=[T]\backslash{\mathcal{T}}_{1}$. Note that ${\mathcal{T}}_{0}$ and ${\mathcal{T}}_{1}$ are sets of iterations such that the model is selected in lines 8 and 13 of Algorithm 4, respectively.
1. Regret incurred in ${\mathcal{T}}_{0}$.
For any $t\in[T]$, let $m_{t}$ denote the stage that model $g_{t}$ is picked at the $t$-th iteration. We have that
$\sum_{t\in{\mathcal{T}}_{0}}(s_{\star}(y_{t})-s_{g_{t}}(y_{t}))\leq$ $\sum_{t\in{\mathcal{T}}_{0}}(\widehat{s}_{g_{\star,t},t}^{m_{t}}(y)-s_{g_{t}}(y_{t}))$ (7)
$\leq$ $\sum_{t\in{\mathcal{T}}_{0}}(\widehat{s}_{g_{t},t}^{m_{t}}(y)-s_{g_{t}}(y_{t}))$
$=$ $\sum_{t\in{\mathcal{T}}_{0}}\left(\widehat{\mu}_{g_{t},t}^{m_{t}}+(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g_{t},t}^{m_{t}}-s_{g_{t}}(y_{t})\right)$
$\leq$ $2(2\eta+\sqrt{\alpha})\sum_{t\in{\mathcal{T}}_{0}}\widehat{\sigma}_{g_{t},t}^{m_{t}}$
$\leq$ $2(2\eta+\sqrt{\alpha})\sum_{t\in{\mathcal{T}}_{0}}T^{-\frac{1}{2}}\leq 2(2\eta+\sqrt{\alpha})\sqrt{T},$
where the second inequality holds by the definition of $g_{t}$ and the fact that $g_{\star,t}\in\widehat{{\mathcal{G}}}^{m_{t}}_{t}$, and the fifth inequality holds by line 7 of Algorithm 4.
2. Regret incurred in ${\mathcal{T}}_{1}$.
$\sum_{t\in{\mathcal{T}}_{1}}(s_{\star}(y_{t})-s_{g_{t}}(y_{t}))=$ $\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}\sum_{t\in\Psi_{g,T+1}^{m}}(s_{\star}(y_{t})-s_{g_{t}}(y_{t}))$ (8)
$\leq$ $\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}2^{3-m}\cdot|\Psi_{g,T+1}^{m}|,$
where the inequality holds by the last statement in Lemma 4. It remains to bound $|\Psi_{g,T+1}^{m}|$. First note that for any $m\in[M]$, we have that
$(2\eta+\sqrt{\alpha})\sum_{t\in\Psi_{g,T+1}^{m}}\widehat{\sigma}_{g,t}^{m}>2^{-m}\cdot|\Psi_{g,T+1}^{m}|$
from line 13 of Algorithm 4. In addition, by a similar statement of (Valko et al.,, 2013, Lemma 4), which is stated in Lemma 10, we have that
$\sum_{t\in\Psi_{g,T+1}^{m}}\widehat{\sigma}_{g,t}^{m}\leq\widetilde{O}\left(\sqrt{d_{\textup{eff}}\left(10+\frac{15}{\alpha}\right)|\Psi_{g,T+1}^{m}|}\right),$ (9)
where $d_{\textup{eff}}$ is defined therein and logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$. Plugging in Equation (8) results in
$\sum_{t\in{\mathcal{T}}_{1}}(s_{\star}(y_{t})-s_{g_{t}}(y_{t}))\leq$ $\widetilde{O}\left((1+\sqrt{\alpha})\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}\sqrt{d_{\textup{eff}}\left(10+\frac{15}{\alpha}\right)|\Psi_{g,T+1}^{m}|}\right)$ (10)
$\leq$ $\widetilde{O}\left((1+\sqrt{\alpha})\sqrt{GM}\sqrt{d_{\textup{eff}}\left(10+\frac{15}{\alpha}\right)\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}|\Psi_{g,T+1}^{m}|}\right)$
$\leq$ $\widetilde{O}\left((1+\sqrt{\alpha})\sqrt{d_{\textup{eff}}\left(10+\frac{15}{\alpha}\right)GT}\right),$
where the second inequality holds by Cauchy-Schwarz inequality.
3. Putting everything together.
Combining Inequalities (7) and (10) leads to
$\textup{Regret}(T)=\left(\sum_{t\in{\mathcal{T}}_{0}}+\sum_{t\in{\mathcal{T}}_{1}}\right)(s_{\star}(y_{t})-s_{g_{t}}(y_{t}))\leq\widetilde{O}\left((1+\sqrt{\alpha})\sqrt{d_{\textup{eff}}\left(10+\frac{15}{\alpha}\right)GT}\right),$
which concludes the proof. ∎
Appendix B Proof in Section 5.2
B.1 Error of KRR estimators with Random Features
Theorem 4.
Assume the error thresholds input to Algorithm 3 satisfy that $\operatorname{\Delta_{\textup{RFF}}},\operatorname{\epsilon_{\textup{RFF}}}\leq 1/2$. Under the same conditions in Lemma 1, with probability at least $1-2\delta$, the quantity $\widetilde{\mu}_{g}$ computed by function Compute_UCB_RFF satisfies that
$|\widetilde{\mu}_{g}-s_{g}(y)|\leq{\mathcal{B}}_{g,1}+(2\eta+\sqrt{\alpha})(\widetilde{\sigma}_{g}+{\mathcal{B}}_{g,2}),$
with the number of features satisfying Inequality (12) and bonus terms ${\mathcal{B}}_{g,1}$ and ${\mathcal{B}}_{g,2}$ given by Equations (11) and (13), where $\eta=\sqrt{2\log(2/\delta)}$. Hence, it holds that $\widetilde{s}_{g}=\widetilde{\mu}_{g}+{\mathcal{B}}_{g,1}+(2\eta+\sqrt{\alpha})(\widetilde{\sigma}_{g}+{\mathcal{B}}_{g,2})\geq s_{g}(y)$.
Proof.
The proof is based on the following two lemmas, which analyze the concentration error of the quantities $\widetilde{\mu}_{g}$ and $\widetilde{\sigma}_{g}$. The detailed proof can be found in Appendix B.3 and B.4, respectively.
Lemma 5 (Concentration of mean using RFF).
Let $\operatorname{\Delta_{\textup{RFF}}},\operatorname{\epsilon_{\textup{RFF}}}\leq 1/2$. Under the same conditions in Lemma 1, with probability at least $1-\delta$, the quantity $\widetilde{\mu}_{g}$ computed by function Compute_UCB_RFF satisfies that
$|\widetilde{\mu}_{g}-\widehat{\mu}_{g}|\leq{\mathcal{B}}_{g,1}:=\alpha^{-1}|\Psi_{g}|\operatorname{\epsilon_{\textup{RFF}}}+\alpha^{-2}\operatorname{\Delta_{\textup{RFF}}}(\|K\|_{2}+\alpha)$ (11)
with the number of features
$s\geq\max\left\{\frac{4(d+2)}{\operatorname{\epsilon_{\textup{RFF}}}^{2}}\log\left(\frac{\sigma_{p}^{2}}{(\delta/2)\cdot\operatorname{\epsilon_{\textup{RFF}}}^{2}}\cdot 2^{8}\right),\frac{8|\Psi_{g}|}{3\alpha}\operatorname{\Delta_{\textup{RFF}}}^{-2}\log\left(\frac{32s_{\alpha}(K)}{\delta}\right)\right\},$ (12)
where $\widehat{\mu}_{g}=(\phi(y))^{\top}\Phi_{g}^{\top}(K+\alpha I)^{-1}v$, and $\sigma_{p}^{2}$ and $s_{\alpha}(\cdot)$ are two quantities defined in Lemmas 8 and 9, respectively.
Lemma 6 (Concentration of variance using RFF).
Let $c>0$ denote a lower bound of $1-\|k_{y}\|_{(K+\alpha I)^{-1}}^{2}$. Then, conditioned on the successful events in Lemma 5, the quantity $\widetilde{\sigma}_{g}$ computed by function Compute_UCB_RFF satisfies that
$|\widetilde{\sigma}_{g}-\widehat{\sigma}_{g}|\leq{\mathcal{B}}_{g,2}:=(c\cdot\alpha)^{-\frac{1}{2}}\left(2|\Psi_{g}|\alpha^{-2}\operatorname{\Delta_{\textup{RFF}}}(\|K\|_{2}+\alpha)+3\alpha^{-1}|\Psi_{g}|\operatorname{\epsilon_{\textup{RFF}}}\right)$ (13)
where $\widehat{\sigma}_{g}=\alpha^{-\frac{1}{2}}\sqrt{k(y,y)-k_{y}^{\top}(K+\alpha I)^{-1}k_{y}}$.
Finally, combining Lemmas 15, and 6, we derive that
$|\widetilde{\mu}_{g}-s_{g}(y)|\leq$ $|\widetilde{\mu}_{g}-\widehat{\mu}_{g}|+\left|\widehat{\mu}_{g}-s_{g}(y)\right|$
$\leq$ ${\mathcal{B}}_{g,1}+(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g}$
$\leq$ ${\mathcal{B}}_{g,1}+(2\eta+\sqrt{\alpha})(\widetilde{\sigma}_{g}+{\mathcal{B}}_{g,2}),$
which concludes the proof. ∎
B.2 Sup-PAK-UCB with Random Fourier Features
Algorithm description.
To apply RFF to Sup-PAK-UCB, we replace function Compute_UCB with Compute_UCB_RFF in Algorithm 4. To achieve the regret bound (6), an important problem is to design (adaptive) error thresholds, i.e., $\operatorname{\epsilon_{\textup{RFF}}}$ and $\operatorname{\Delta_{\textup{RFF}}}$, when computing UCB at each stage $m$ and iteration $t$. We prove the regret bound in the following theorem.
Theorem 5 (Regret of Sup-RFF-UCB).
Under Assumption 1, with probability at least $1-\delta$, Sup-RFF-UCB attains the regret bound (6), where $\eta=\sqrt{2\log(4(\log T)GT/\delta)}$ and sequence of error thresholds input to function Compute_UCB_RFF satisfying Equation (14).
Proof.
The proof is similar to the proof of Theorem 3. First, combining Lemmas 8 and 3, the following lemma can be proved by the exact same analysis for Lemma 4.
Lemma 7.
With probability at least $1-(2MGT)\delta$, for any iteration $t\in[T]$ and stage $m\in[M]$, the following hold:
  • $|\widetilde{\mu}_{g,t}^{m}-s_{g}(y_{t})|\leq{\mathcal{B}}_{g,1,t}^{m}+(2\eta+\sqrt{\alpha})(\widetilde{\sigma}_{g,t}^{m}+{\mathcal{B}}_{g,2,t}^{m})$ for any $g\in\widehat{{\mathcal{G}}}^{m}_{t}$,
  • $\operatorname*{arg\,max}_{g\in{\mathcal{G}}}s_{g}(y_{t})\in\widehat{{\mathcal{G}}}^{m}_{t}$, and
  • $s_{\star}(y_{t})-s_{g}(y_{t})\leq 2^{3-m}$ for any $g\in\widehat{{\mathcal{G}}}^{m}_{t}$.
where the first statement is guaranteed by Theorem 4, ${\mathcal{B}}_{g,1,t}^{m}$ and ${\mathcal{B}}_{g,2,t}^{m}$ are the bonus (11) and (13) computed at the $m$-th stage of iteration $t$.
Next, for iterations in ${\mathcal{T}}_{0}$ (model $g_{t}$ is picked in line 8 of Algorithm 4), we still have
$\sum_{t\in{\mathcal{T}}_{0}}(s_{\star}(y_{t})-s_{g_{t}}(y_{t}))\leq\widetilde{O}(\sqrt{\alpha T}).$
Further, for iterations in ${\mathcal{T}}_{1}$ (model $g_{t}$ is picked in line 13 of Algorithm 4), the third statement in the above lemma and line 14 of Algorithm 3 ensure that
$\sum_{t\in{\mathcal{T}}_{1}}(s_{\star}(y_{t})-s_{g_{t}}(y_{t}))$
$\leq$ $\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}2^{3-m}\cdot|\Psi_{g,T+1}^{m}|$
$<$ $8\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}\sum_{t\in\Psi_{g,T+1}^{m}}\left({\mathcal{B}}_{g_{t},1,t}^{m_{t}}+(2\eta+\sqrt{\alpha})\left(\widetilde{\sigma}_{g,t}^{m}+{\mathcal{B}}_{g_{t},2,t}^{m_{t}}\right)\right)$
$\leq$ $8\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}\sum_{t\in\Psi_{g,T+1}^{m}}\left({\mathcal{B}}_{g_{t},1,t}^{m_{t}}+(2\eta+\sqrt{\alpha})\left(\widehat{\sigma}_{g,t}^{m}+2{\mathcal{B}}_{g_{t},2,t}^{m_{t}}\right)\right)$
$=$ $8\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}\left(\sum_{t\in\Psi_{g,T+1}^{m}}\left({\mathcal{B}}_{g_{t},1,t}^{m_{t}}+2(2\eta+\sqrt{\alpha}){\mathcal{B}}_{g_{t},2,t}^{m_{t}}\right)+(2\eta+\sqrt{\alpha})\sum_{t\in\Psi_{g,T+1}^{m}}\widehat{\sigma}_{g,t}^{m}\right).$
Note that the upper bound of the second term has been derived in Equation (9). It remains to bound the first term. Essentially, we will find a sequence of error thresholds, and hence the number of features defined in Inequality (12), such that the first term is bounded by $\widetilde{O}(\sqrt{GT})$. For convenience, we introduce the following notations:
Additional notations. For any iteration $t\in[T]$, model $g\in{\mathcal{G}}$, and stage $m\in[M]$, we define $K_{g,t}^{m}:=\Phi_{g,t}^{m}(\Phi_{g,t}^{m})^{\top}$, where $\Phi_{g,t}^{m}:=[\phi(y_{i})^{\top}]_{i\in\Psi_{g,t}^{m}}$. In addition, we denote by $0<c_{g,t}^{m}\leq 1$ the lower bound in Lemma 6 corresponding to $y_{t}$ and $K_{g,t}^{m}$. Let
$\epsilon_{\textup{RFF},g,t}^{m}\leq t^{-\frac{1}{2}}(|\Psi_{g,t}^{m}|)^{-1}\sqrt{G\cdot c_{g,t}^{m}},\quad\Delta_{\textup{RFF},g,t}^{m}\leq t^{-\frac{1}{2}}(|\Psi_{g,t}^{m}|(\|K_{g,t}^{m}\|_{2}+\alpha))^{-1}\sqrt{G\cdot c_{g,t}^{m}}$ (14)
denote the (upper bound of) error thresholds input to function Compute_UCB_RFF.
$\sum_{g\in{\mathcal{G}},m\in[M]}\sum_{t\in\Psi_{g,T+1}^{m}}{\mathcal{B}}_{g_{t},1,t}^{m}$ (15)
$=$ $\sum_{g\in{\mathcal{G}},m\in[M]}\sum_{t\in\Psi_{g,T+1}^{m}}\left(\alpha^{-1}|\Psi_{g,t}^{m}|\epsilon_{\textup{RFF},g,t}^{m}+\alpha^{-2}\Delta_{\textup{RFF},g,t}^{m}(\|K_{g,t}^{m}\|_{2}+\alpha)\right)$
$\leq$ $\sqrt{G}\sum_{g\in{\mathcal{G}},m\in[M]}\sum_{t\in\Psi_{g,T+1}^{m}}(\alpha^{-1}t^{-\frac{1}{2}}+\alpha^{-2}t^{-\frac{1}{2}})$
$\leq$ $\sqrt{G}\sum_{t=1}^{T}(\alpha^{-1}t^{-\frac{1}{2}}+\alpha^{-2}t^{-\frac{1}{2}})$
$\leq$ $O\left((\alpha^{-1}+\alpha^{-2})\sqrt{GT}\right)$
where the first inequality holds by the fact that each $t\in[T]$ appears in at most one index set.
$\sum_{g\in{\mathcal{G}},m\in[M]}\sum_{t\in\Psi_{g,T+1}^{m}}{\mathcal{B}}_{g_{t},2,t}^{m}$ (16)
$\leq$ $\sum_{g\in{\mathcal{G}},m\in[M]}\sum_{t\in\Psi_{g,T+1}^{m}}(c\cdot\alpha)^{-\frac{1}{2}}\left(2|\Psi_{g}|\alpha^{-2}\Delta_{\textup{RFF},g,t}^{m}(\|K_{g,t}^{m}\|_{2}+\alpha)+3\alpha^{-1}|\Psi_{g,t}^{m}|\epsilon_{\textup{RFF},g,t}^{m}\right)$
$\leq$ $\sqrt{G}\sum_{g\in{\mathcal{G}},m\in[M]}\sum_{t\in\Psi_{g,T+1}^{m}}\left(2\alpha^{-\frac{5}{2}}t^{-\frac{1}{2}}+3\alpha^{-\frac{3}{2}}t^{-\frac{1}{2}}\right)$
$\leq$ $\sqrt{G}\sum_{t=1}^{T}\left(2\alpha^{-\frac{5}{2}}t^{-\frac{1}{2}}+3\alpha^{-\frac{3}{2}}t^{-\frac{1}{2}}\right)$
$\leq$ $O\left((\alpha^{-\frac{5}{2}}+\alpha^{-\frac{3}{2}})\sqrt{GT}\right)$
Therefore, we conclude the proof. ∎
B.3 Proof of Lemma 5
Proof.
For convenience, we define $\tilde{k}_{y}:=\widetilde{\Phi}_{g}(\varphi(y))\in{\mathbb{R}}^{|\Psi_{g}|},Q:=(K+\alpha I)^{-1}\in{\mathbb{R}}^{|\Psi_{g}|\times|\Psi_{g}|}$, and $\widetilde{Q}:=(\widetilde{K}+\alpha I)^{-1}\in{\mathbb{R}}^{|\Psi_{g}|\times|\Psi_{g}|}$, where $\widetilde{K}:=\widetilde{\Phi}_{g}\widetilde{\Phi}_{g}^{\top}$. Using the same notations in the proof of Lemma 1, we obtain that
$|\widetilde{\mu}_{g}-\widehat{\mu}_{g}|=$ $\left|(\varphi(y))^{\top}\widetilde{\Phi}_{g}^{\top}(\widetilde{K}+\alpha I)^{-1}v-k_{y}^{\top}(K+\alpha I)^{-1}v\right|$ (17)
$=$ $\left|\tilde{k}_{y}^{\top}\widetilde{Q}v-k_{y}^{\top}Qv\right|$
$\leq$ $\left|\tilde{k}_{y}^{\top}(\widetilde{Q}-Q)v\right|+\left|(\tilde{k}_{y}-k_{y})^{\top}Qv\right|,$
where we use Equation (22) to derive $\widetilde{\mu}_{g}=(\varphi(y))^{\top}(\widetilde{\Phi}_{g}^{\top}\widetilde{\Phi}_{g}+\alpha I)^{-1}\widetilde{\Phi}_{g}^{\top}v=(\varphi(y))^{\top}\widetilde{\Phi}_{g}^{\top}(\widetilde{K}+\alpha I)^{-1}v$ in the first equation.
1. Bounding $|(\tilde{k}_{y}-k_{y})^{\top}Qv|$.
We evoke (Rahimi and Recht, 2007b, , Claim 1), which is rewritten in Lemma 8 using our notations. For a desired threshold $\operatorname{\epsilon_{\textup{RFF}}}>0$, set
$s=\frac{4(d+2)}{\operatorname{\epsilon_{\textup{RFF}}}^{2}}\log\left(\frac{\sigma_{p}^{2}}{(\delta/2)\cdot\operatorname{\epsilon_{\textup{RFF}}}^{2}}\cdot 2^{8}\right).$
Then, with probability at least $1-\frac{\delta}{2}$, it holds that $\sup_{y,y^{\prime}\in{\mathcal{Y}}}|(\varphi(y))^{\top}\varphi(y^{\prime})-k(y,y^{\prime})|\leq\operatorname{\epsilon_{\textup{RFF}}}$, and hence $\|\tilde{k}_{y}-k_{y}\|_{\infty}\leq\operatorname{\epsilon_{\textup{RFF}}}$. Therefore, we obtain
$|(\tilde{k}_{y}-k_{y})^{\top}Qv|\leq\|\tilde{k}_{y}-k_{y}\|_{2}\cdot\|Q\|_{2}\cdot\|v\|_{2}\leq\operatorname{\epsilon_{\textup{RFF}}}\sqrt{|\Psi_{g}|}\cdot\alpha^{-1}\cdot\sqrt{|\Psi_{g}|}=\alpha^{-1}|\Psi_{g}|\operatorname{\epsilon_{\textup{RFF}}},$ (18)
where the last inequality holds by $\|Q\|_{2}=\lambda_{\min}^{-1}(K+\alpha I)\leq\alpha^{-1}$ and $\|v\|_{\infty}\leq 1$.
2. Bounding $|\tilde{k}_{y}^{\top}(\widetilde{Q}-Q)v|$.
Note that
$|\tilde{k}_{y}^{\top}(\widetilde{Q}-Q)v|\leq$ $\|\tilde{k}_{y}\|_{2}\cdot\|\widetilde{Q}-Q\|_{2}\cdot\|v\|_{2}$
$\leq$ $\sqrt{2|\Psi_{g}|}\cdot\|\widetilde{Q}-Q\|_{2}\cdot\sqrt{|\Psi_{g}|}$
where the first inequality holds by the fact that $(\varphi(y))^{\top}\varphi(y_{i})=(2/s)\sum_{j=1}^{s}\cos(w_{j}^{\top}y+b_{j})\cos(w_{i,j}^{\top}+b_{i,j})\leq 2$. To bound $\|\widetilde{Q}-Q\|_{2}$, we evoke (Avron et al.,, 2017, Theorem 7), which is rewritten in Lemma 9. For a desired threshold $\operatorname{\Delta_{\textup{RFF}}}\leq 1/2$, the following inequality holds with probability at least $1-\frac{\delta}{2}$:
$(1-\operatorname{\Delta_{\textup{RFF}}})(K+\alpha I)\preceq\widetilde{K}+\alpha I$
for $s\geq\frac{8|\Psi_{g}|}{3\alpha}\operatorname{\Delta_{\textup{RFF}}}^{-2}\log(32s_{\alpha}(K)/\delta)$. By Sherman-Morrison-Woodbury formula, i.e., $A^{-1}-B^{-1}=A^{-1}(B-A)B^{-1}$ where $A$ and $B$ are invertible, we derive
$\|(\widetilde{K}+\alpha I)^{-1}-(K+\alpha I)^{-1}\|_{2}$ (19)
$\leq$ $\|(\widetilde{K}+\alpha I)^{-1}\|_{2}\cdot\|(K+\alpha I)-(\widetilde{K}+\alpha I)\|_{2}\cdot\|(K+\alpha I)^{-1}\|_{2}$
$\leq$ $\alpha^{-2}\operatorname{\Delta_{\textup{RFF}}}(\|K\|_{2}+\alpha)$
where the last inequality holds by the fact that $\|(\widetilde{K}+\alpha I)^{-1}\|_{2},\|(K+\alpha I)^{-1}\|_{2}\leq\alpha^{-1}$ and $\|(K+\alpha I)-(\widetilde{K}+\alpha I)\|_{2}\leq\|\operatorname{\Delta_{\textup{RFF}}}(K+\alpha I)\|_{2}\leq\operatorname{\Delta_{\textup{RFF}}}(\|K\|_{2}+\alpha)$.
3. Putting everything together.
Combining Equations (18) and (LABEL:eq43), with probability at least $1-\delta$, it holds that
$|\widetilde{\mu}_{g}-\widehat{\mu}_{g}|\leq\alpha^{-1}|\Psi_{g}|\operatorname{\epsilon_{\textup{RFF}}}+\alpha^{-2}\operatorname{\Delta_{\textup{RFF}}}(\|K\|_{2}+\alpha)$
when
$s\geq\max\left\{\frac{4(d+2)}{\operatorname{\epsilon_{\textup{RFF}}}^{2}}\log\left(\frac{\sigma_{p}^{2}}{(\delta/2)\cdot\operatorname{\epsilon_{\textup{RFF}}}^{2}}\cdot 2^{8}\right),\frac{8|\Psi_{g}|}{3\alpha}\operatorname{\Delta_{\textup{RFF}}}^{-2}\log\left(\frac{32s_{\alpha}(K)}{\delta}\right)\right\},$
which concludes the proof. ∎
B.4 Proof of Lemma 6
Proof.
We use the same notations in the proof of Lemma 5. Let $c$ denote a lower bound of $1-\|k_{y}\|_{(K+\alpha I)^{-1}}^{2}$. Note that
$|\widetilde{\sigma}_{g}-\widehat{\sigma}_{g}|$ (20)
$=$ $\alpha^{-\frac{1}{2}}\left|\sqrt{1-(\varphi(y))^{\top}\widetilde{\Phi}_{g}^{\top}(\widetilde{K}+\alpha I)^{-1}\widetilde{\Phi}_{g}(\varphi(y))}-\sqrt{1-k_{y}^{\top}(K+\alpha I)^{-1}k_{y}}\right|$
$\leq$ $(c\cdot\alpha)^{-\frac{1}{2}}\left|(\varphi(y))^{\top}\widetilde{\Phi}_{g}^{\top}(\widetilde{K}+\alpha I)^{-1}\widetilde{\Phi}_{g}(\varphi(y))-k_{y}^{\top}(K+\alpha I)^{-1}k_{y}\right|$
$=$ $(c\cdot\alpha)^{-\frac{1}{2}}\left|\tilde{k}_{y}^{\top}\widetilde{Q}\tilde{k}_{y}-k_{y}^{\top}Qk_{y}\right|$
$\leq$ $(c\cdot\alpha)^{-\frac{1}{2}}\left(\left|\tilde{k}_{y}^{\top}(\widetilde{Q}-Q)\tilde{k}_{y}\right|+\left|(\tilde{k}_{y}-k_{y})^{\top}Q\tilde{k}_{y}\right|+\left|k_{y}^{\top}Q(\tilde{k}_{y}-k_{y})\right|\right)$
$\leq$ $(c\cdot\alpha)^{-\frac{1}{2}}\left(\|\tilde{k}_{y}\|_{2}^{2}\|\widetilde{Q}-Q\|_{2}+\|\tilde{k}_{y}-k_{y}\|_{2}\|\tilde{k}_{y}\|_{2}\|Q\|_{2}+\|\tilde{k}_{y}-k_{y}\|_{2}\|k_{y}\|_{2}\|Q\|_{2}\right)$
$\leq$ $(c\cdot\alpha)^{-\frac{1}{2}}\left(2|\Psi_{g}|\alpha^{-2}\operatorname{\Delta_{\textup{RFF}}}(\|K\|_{2}+\alpha)+3\alpha^{-1}|\Psi_{g}|\operatorname{\epsilon_{\textup{RFF}}}\right)$
which concludes the proof. ∎
B.5 Analysis of Lemma 2
Proof.
Solving KRR with $n$ regression data requires $\Theta(n^{3})$ time and $\Theta(n^{2})$ space. Hence, by the convexity of the cubic and quadratic functions, the time for Compute_UCB scales with $\Theta(\sum_{g\in{\mathcal{G}}}n_{g}^{3})=O(t^{3}/G^{2})$, and the space scales with $\Theta(\sum_{g\in{\mathcal{G}}}n_{g}^{2})=O(t^{2}/G)$, where $n_{g}:=|\Psi_{g}|$ is the visitation to any model $g\in{\mathcal{G}}$ up to iteration $t$, and we have $\sum_{g\in{\mathcal{G}}}n_{g}=t$. On the other hand, solving KRR with $n$ regression data and random features of size $s$ requires $O(ns^{2})$ time and $O(ns)$ space. Therefore, the time for Compute_UCB_RFF scales with $O(\sum_{g\in{\mathcal{G}}}n_{g}s^{2})=O(ts^{2})$, and the space scales with $O(\sum_{g\in{\mathcal{G}}}n_{g}s)=O(ts)$. ∎
Appendix C Auxiliary Lemmas
C.1 Proof of Lemma 1
Proof.
We rewrite the proof using the notations in Section 5. Obviously, Equation (5) holds when the index set $\Psi_{g}$ is empty. In the following, we consider non-empty $\Psi_{g}$. Let $\Phi_{g}:=[\phi(y_{i})^{\top}]_{i\in\Psi_{g}}$. Note that $k_{y}=[k(y,y_{i})]^{\top}_{i\in\Psi_{g}}=\Phi_{g}(\phi(y))$ and $K=[k(y_{i},y_{j})]_{i,j\in\Psi_{g}}=\Phi_{g}\Phi_{g}^{\top}$. We have
$\widehat{\mu}_{g}-s_{g}(y)=$ $(\phi(y))^{\top}\Phi_{g}^{\top}(K+\alpha I)^{-1}v-(\phi(y))^{\top}w_{g}^{\star}$ (21)
$=$ $(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}\Phi_{g}^{\top}v-(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)w_{g}^{\star}$
$=$ $(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}\Phi_{g}^{\top}(v-\Phi_{g}w_{g}^{\star})-\alpha(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}w_{g}^{\star},$
where the second equation holds by the positive definiteness of both matrices $(K+\alpha I)$ and $(\Phi_{g}^{\top}\Phi_{g}+\alpha I)$ and hence
$\Phi_{g}^{\top}(K+\alpha I)^{-1}=(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}\Phi_{g}^{\top}.$ (22)
1. Bounding $(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}\Phi_{g}^{\top}(v-\Phi_{g}w_{g}^{\star})$.
Note that the scores $\{s_{t}:t\in\Psi_{g}\}$ are independent by the construction of $\Phi_{g}$ and $\mathbb{E}[s_{t}]=(w_{g}^{\star})^{\top}\phi(y_{t})$, we have that
$(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}\Phi_{g}^{\top}(v-\Phi_{g}w_{g}^{\star})=\sum_{i=1}^{|\Psi_{g}|}[(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}\Phi_{g}^{\top}]_{i}\cdot[v-\Phi_{g}w_{g}^{\star}]_{i}$
are the summation of zero mean independent random variables, where we denote by $[\cdot]_{i}$ the $i$-th element of a vector. Further, each variable satisfies that
$\left|[(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}\Phi_{g}^{\top}]_{i}\cdot[v-\Phi_{g}w_{g}^{\star}]_{i}\right|$
$\leq$ $\|(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}\Phi_{g}^{\top}\|\cdot|[v-\Phi_{g}w_{g}^{\star}]_{i}|$
$\leq$ $\sqrt{(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}\Phi_{g}^{\top}\Phi_{g}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}(\phi(y))}\cdot(1+\|w_{g}^{\star}\|)$
$\leq$ $2\widehat{\sigma}_{g}$
where the last inequality holds by $\|w_{g}^{\star}\|\leq 1$ and the second inequality holds by
$\widehat{\sigma}_{g}=$ $\alpha^{-\frac{1}{2}}\sqrt{k(y,y)-k_{y}^{\top}(K+\alpha I)^{-1}k_{y}}$
$=$ $\alpha^{-\frac{1}{2}}\sqrt{(\phi(y))^{\top}(\phi(y))-(\phi(y))^{\top}\Phi_{g}^{\top}(K+\alpha I)^{-1}\Phi_{g}(\phi(y))}$
$=$ $\alpha^{-\frac{1}{2}}\sqrt{(\phi(y))^{\top}\left(I-(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}\Phi_{g}^{\top}\Phi_{g}\right)(\phi(y))}$
$=$ $\sqrt{(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}(\phi(y))},$
Then, by Azuma-Hoeffding inequality, it holds that
${\mathbb{P}}\left(|(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}\Phi_{g}^{\top}(v-\Phi_{g}w_{g}^{\star})|>2\eta\widehat{\sigma}_{g}\right)$ (23)
$\leq$ $2\exp\left(-\frac{\widehat{\sigma}_{g}^{2}\eta^{2}}{2|\Psi_{g}|\widehat{\sigma}_{g}^{2}}\right)$
$\leq$ $2\exp(-\eta^{2}/2)$
2. Bounding $\alpha(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}w_{g}^{\star}$.
By the Cauchy-Schwarz inequality, it holds that
$\left|(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}w_{g}^{\star}\right|$ (24)
$\leq$ $\|w_{g}^{\star}\|\cdot\|(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}\|$
$=$ $\|w_{g}^{\star}\|\cdot\sqrt{(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}\alpha^{-1}\alpha I(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}(\phi(y))}$
$\leq$ $\alpha^{-1/2}\sqrt{(\phi(y))^{\top}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}(\Phi_{g}^{\top}\Phi_{g}+\alpha I)(\Phi_{g}^{\top}\Phi_{g}+\alpha I)^{-1}(\phi(y))}$
$=$ $\alpha^{-1/2}\widehat{\sigma}_{g},$
where the second inequality holds by the positive definiteness of $\Phi_{g}^{\top}\Phi_{g}$.
3. Putting everything together.
Plugging (23) and (24) in (21) and setting $\delta=2\exp(-\eta^{2}/2)$ concludes the proof. ∎
C.2 Proof of Lemma 4
Proof.
The first statement holds by both Lemma 3 and Lemma 1, and a uniform bound over all $t\in[T],g\in{\mathcal{G}}$, and $m\in[M]$. Let $g_{\star,t}:=\operatorname*{arg\,max}_{g\in{\mathcal{G}}}s_{g}(y_{t})$ is the optimal model for prompt $y_{t}$ and $\widehat{g}_{\star,t}^{m}:=\operatorname*{arg\,max}_{g\in\widehat{{\mathcal{G}}}^{m}_{t}}\widehat{s}_{g,t}^{m}$ is optimistic model at stage $m$.
To show the second statement, first note that $g_{\star,t}\in\widehat{{\mathcal{G}}}^{1}_{t}$. Assume $g_{\star,t}\in\widehat{{\mathcal{G}}}^{m}_{t}$ for some $m\in[M-1]$. Then, by the first statement, we obtain that $\widehat{s}_{g_{\star,t},t}^{m}-\max_{g\in\widehat{{\mathcal{G}}}^{m}_{t}}\widehat{s}_{g,t}^{m}\geq s_{g_{\star,t}}(y_{t})-(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{g_{\star,t}}^{m}-(s_{\widehat{g}_{\star,t}^{m}}(y_{t})+(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{\widehat{g}_{\star,t}^{m},t}^{m})\geq-(2\eta+\sqrt{\alpha})\cdot(\widehat{\sigma}_{g_{\star,t}}^{m}+\widehat{\sigma}_{\widehat{g}_{\star,t}^{m},t}^{m})\geq 2\cdot 2^{-m}=2^{1-m}$, which ensures $g_{\star,t}\in\widehat{{\mathcal{G}}}^{m+1}_{t}$.
Finally, by the first two statements, we have that $s_{\star}(y_{t})-s_{g}(y_{t})\leq\widehat{s}_{g_{\star,t},t}^{m}+2(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{g_{\star,t}}^{m}-(\widehat{s}_{g,t}^{m}-2(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{g}^{m,t})\leq 2\cdot 2^{1-m}=2^{3-m}$. We conclude the proof. ∎
C.3 Useful Lemmas
Lemma 8 (Rahimi and Recht, 2007b , Claim 1).
Let ${\mathcal{M}}$ be a compact subset of ${\mathbb{R}}^{d}$ with diameter $\textup{diam}({\mathcal{M}})$. Then, for the mapping $\varphi:{\mathbb{R}}^{d}\to{\mathbb{R}}^{s}$ defined in Equation (3), we have
${\mathbb{P}}\left(\sup_{y,y^{\prime}\in{\mathcal{M}}}|(\varphi(y))^{\top}\varphi(y^{\prime})-k(y,y^{\prime})|\geq\epsilon\right)\leq 2^{8}\left(\frac{\sigma_{p}\cdot\textup{diam}({\mathcal{M}})}{\epsilon}\right)^{2}\exp\left(-\frac{s\epsilon^{2}}{4(d+2)}\right),$
where $\sigma_{p}^{2}:=\mathbb{E}_{p}[\omega^{\top}\omega]$ is the second moment of the Fourier transform of $k$ [^6] For the RBF kernel with parameter $\sigma^{2}$, i.e., $k_{\textup{RBF}}(y,y^{\prime})=\exp(-\frac{1}{2\sigma^{2}}\|y-y^{\prime}\|_{2}^{2})$, we have $\sigma_{p_{\textup{RBF}}}^{2}=\frac{d}{\sigma^{2}}$. Further, $\sup_{y,y^{\prime}\in{\mathcal{M}}}|(\varphi(y))^{\top}\varphi(y^{\prime})-k(y,y^{\prime})|\leq\epsilon$ with any constant probability when $s=\Omega((d/\epsilon^{2})\log(\sigma_{p}\cdot\textup{diam}({\mathcal{M}})/\epsilon))$.
Lemma 9 (Avron et al., (2017), Theorem 7).
Let $K=[k(y_{i},y_{j})]_{i,j\in[n]}$ denote the Gram matrix of $\{y_{i}\in{\mathbb{R}}^{d}\}_{i=1}^{n}$, where $k$ is a shift-invariant kernel function. Let $\Delta\leq 1/2$ and $\delta\in(0,1)$. Assume that $\|K\|_{2}\geq\alpha$. If we use $s\geq\frac{8n}{3\alpha}\Delta^{-2}\log(16s_{\alpha}(K)/\delta)$ random Fourier features, then with probability at least $1-\delta$, it holds that
$(1-\Delta)(K+\alpha I)\preceq\widetilde{K}+\alpha I\preceq(1+\Delta)(K+\alpha I)$
where $s_{\alpha}(K):=\textup{Tr}[(K+\alpha I)^{-1}K]$ and we denote by $\widetilde{K}=[(\varphi(y_{i}))^{\top}(\varphi(y_{j}))]_{i,j\in[n]}$ the approximated Gram matrix using $s\in{\mathbb{N}}_{+}$ random Fourier features, where $\varphi:{\mathbb{R}}^{d}\to{\mathbb{R}}^{s}$ is the feature mapping.
Lemma 10 (Valko et al., (2013), Lemma 4).
For any model $g\in{\mathcal{G}}$ and stage $m\in[M]$, let $\lambda_{g,1}^{m}\geq\lambda_{g,2}^{m}\geq\cdots$ denote the eigenvalues (in the decreasing order) of the matrix $(\Phi_{g}^{m})^{\top}\Phi_{g}^{m}+\alpha I$, where $\Phi_{g}^{m}=[\phi(y_{i})^{\top}]_{i\in\Psi_{g,T+1}^{m}}$. Then, for any iteration $t\in[T]$, it holds that
$\sum_{t\in\Psi_{g,T+1}^{m}}\widehat{\sigma}_{g,t}^{m}\leq\widetilde{O}\left(\sqrt{d_{\textup{eff}}\left(10+\frac{15}{\alpha}\right)|\Psi_{g,T+1}^{m}|}\right)$
where $d_{\textup{eff}}:=\max_{g\in{\mathcal{G}},m\in[M]}\min\{j\in{\mathbb{N}}_{+}:j\alpha\log T\geq\Lambda_{g,j}^{m}\}$ and $\Lambda_{g,j}^{m}:=\sum_{i>j}\lambda_{g,i}^{m}-\alpha$ is the effective dimension.
Appendix D Additional Experimental Details and Results
1. Varying rankings of text-to-image models.
We provide more examples showing that prompt-based generative models can outperform for text prompts from certain categories while underperforming for other text categories (see Figures 3, 4, and 6).
Stable Diffusion v1.5 PixArt-$\bm{\alpha}$ Examples (clockwise)
Prompts of Type “dog”
Refer to caption Refer to caption 1. “a woman sitting with     her dog and a toy” 2. “the little dog stands     near the toilet in     a small bathroom” 3. “a dog on a leash     drinking water     from a water bottle” 4. “a dog laying on     the floor next to a door”
avg.CLIPScore $36.37\ (\pm 0.13)$ $\bm{37.24}\ (\pm 0.09)$
Prompts of Type “car”
Refer to caption Refer to caption 1. “a motorcycle is on     the road behind a car” 2. “two cars and a motorcycle     on a road being crossed     by a herd of elephants” 3. “a car that had run     over a red fire hydrant” 4. “a taxi driving down     a city street below     tall white buildings”
avg.CLIPScore $\bm{36.10}\ (\pm 0.06)$ $35.68\ (\pm 0.15)$
Figure 3: Prompt-based generated images from Stable Diffusion v1.5 and PixArt-$\alpha$: Stable Diffusion v1.5 attains a higher CLIPScore in generating type-2 prompts (36.10 versus 35.68) while underperforms for type-1 prompts (36.37 versus 37.24).
UniDiffuser PixArt-$\bm{\alpha}$ Examples (clockwise)
Prompts of Type “train”
Refer to caption Refer to caption 1. “the rusted out remains     of a small railway line” 2. “a skier stands in the     snow outside of a train 3. “train cars parked on a     train track near a pile     of construction material” 4. “several people on horses     with a train car     in the background”
avg.CLIPScore $\bm{35.29}\ (\pm 0.08)$ $34.25\ (\pm 0.12)$
Prompts of Type “baseball bat”
Refer to caption Refer to caption 1. “a man in red shirt     holding a baseball bat” 2. “a woman holding a     baseball bat with her     head resting on it” 3. “baseball player in the     batter’s box hitting     a baseball 4. “hind view of a baseball     player, an umpire, and     a catcher”
avg.CLIPScore $32.51\ (\pm 0.05)$ $\bm{34.30}\ (\pm 0.04)$
Figure 4: Prompt-based generated images from UniDiffuser (Bao et al.,, 2023) and PixArt-$\alpha$: UniDiffuser attains a higher CLIPScore in generating type “train” prompts (35.29 versus 34.25) while underperforms for type “baseball bat” prompts (32.51 versus 34.30).
UniDiffuser Stable Diffusion v1.5 Examples (clockwise)
Prompts of Type “elephant”
Refer to caption Refer to caption 1. “an elephant is carrying     people across a     forested area” 2. “an elephant has a     sheet on its back 3. “a small gray elephant     standing on a beach     next to a lake” 4. “a large elephant in an     open field approaching     a vehicle”
avg.CLIPScore $\bm{36.67}\ (\pm 0.05)$ $35.08\ (\pm 0.06)$
Prompts of Type “fire hydrant”
Refer to caption Refer to caption 1. “a fire hydrant in a clump     of flowering bushes” 2. “a fire hydrant on a gravel     ground with a fence     behind it” 3. “a fire hydrant that     is in the grass” 4. “a toy Ford truck next to     a fire hydrant”
avg.CLIPScore $35.11\ (\pm 0.14)$ $\bm{37.23}\ (\pm 0.05)$
Figure 6: Prompt-based generated images from UniDiffuser and Stable Diffusion v1.5: UniDiffuser attains a higher CLIPScore in generating type “elephant” prompts (36.67 versus 35.08) while underperforms for type “fire hydrant” prompts (35.11 versus 37.23).
2. Prompt-based selection over three T2I models.
See Figure LABEL:fig:t2i4pm.
3. Adaptation to new models and prompts.
See Figures LABEL:fig:varying_model and LABEL:fig:varying_prompt.
4. Implementation details.
We use the CLIP-based features of the prompts as the context vector (Cherti et al.,, 2023) for tasks of T2I and T2V generation, and we use the CLIP-based features of the images in the task of image captioning. We set both the exploration and regularization parameters $\alpha,\eta=1$ in all the experiments. Two hyperparameters have to be chosen. The first one is the parameter $\gamma$ in the polynomial and radial basis function (RBF) kernels, which are given by
$k_{\textup{poly3}}^{\gamma}(x_{1},x_{2})=(\gamma\cdot x_{1}^{\top}x_{2}+1)^{3},\quad k_{\textup{RBF}}^{\gamma}(x_{1},x_{2})=\exp(-\gamma\cdot\|x_{1}-x_{2}\|^{2}).$
In the experiments, we select $\gamma$ to be 5 and 1 for the polynomial and RBF kernel functions, respectively. The second hyperparameter is the number of random features in the RFF-UCB algorithm. In addition, the features are generated once for each sample and stored to save the computation of the RFF-UCB algorithm.
5. Ablation study on hyperparameters.
[…]
6. Comparison of running time.
We compare the running time of PAK-UCB and RFF-UCB, and the results are summarized in Figure 8.
Figure 8: Running time: The execution time of PAK-UCB-RBF (PAK-UCB using the RBF kernel) and RFF-UCB on Setup 4. PAK-UCB-RBF takes around 10 minutes to finish 2,000 iterations of model selection, while RFF-UCB uses less than 2 minutes. Results are averaged over 20 trials.
7. Results on Other Conditional Generation Tasks
1. Text-to-Image (T2I). In this setup, we synthesize five T2I generators based on Stable Diffusion 2  [^7] https://huggingface.co/stabilityai/stable-diffusion-2, where each generator is an “expert” in generating images corresponding to a prompt type. The prompts are captions in the MS-COCO dataset from five categories: dog, car, carrot, cake, and bowl. At each iteration, a caption is drawn from a (random) category, and an image is generated from Stable Diffusion 2. If the learner does not select the expert generator, then we add Gaussian noise to the generated image. Examples are visualized in Figure 9.
Generator 1 2 3 4 5
Type 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Type 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Type 3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Type 4 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Type 5 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Generated images with noise perturbations: Each row and column display the generated images from a synthetic generator according to one single type of prompts. Images generated by the expert models are framed by green boxes. Gaussian noises are applied to non-expert models.
2. Image Captioning. In this setup, the images are chosen from the MS-COCO dataset from five categories: dog, car, carrot, cake, and bowl. We synthesize five expert generators based on the vit-gpt2 model in the Transformers repository. [^8] https://huggingface.co/nlpconnect/vit-gpt2-image-captioning If a non-expert generator is chosen, then the caption is generated from the noisy image perturbed by Gaussian noises. Examples are visualized in Figure 11. The numerical results are summarized in Figure LABEL:fig:imgcap.
Example 1 Example 2 Example 3
Clean
Refer to caption “a person on a surfboard in the air above the water” Refer to caption “a man standing in a kitchen with a dog” Refer to caption “a bowl filled with ice cream and strawberries
28.57 40.53 27.98
Noisy
Refer to caption “a blurry photo of a skateboarder flying through the air” Refer to caption “a cat that is standing in the grass” Refer to caption “a blue and white bowl filled with water”
24.72 13.27 25.83
Figure 11: Generated captions for the clean and noise-perturbed images from vit-gpt2 and the corresponding CLIPScore.
3. Text-to-Video (T2V). We provide numerical results on a synthetic T2V setting. Specifically, both the captions and videos are randomly selected from the following five categories of the MSR-VTT dataset (Xu et al.,, 2016): sports/action, movie/comedy, vehicles/autos, music, and food/drink. Each of the five synthetic arms corresponds to an expert in “generating” videos from a single category. Gaussian noises are applied to the video for non-experts. The results are summarized in Figure LABEL:fig:t2v.